OpenStudy (sugarmochi):

A trundle wheel is to have a circumference of 1 metre. What must be its diameter, to the nearest millimeter

OpenStudy (mrnood):

You now know how to do this C= pi *D

OpenStudy (sugarmochi):

but how do you do it the other way around?

OpenStudy (sugarmochi):

i got stuck on how to do it back wards

OpenStudy (mrnood):

if oyu have an equation A=B*C and you want toknow what C is Divide both sides by B \[\frac{ A }{ B }=\frac{ BC }{ B } =C\]

OpenStudy (mrnood):

In this case - divide both sides by pi

OpenStudy (sugarmochi):

so pi÷1m? then what is the other side?

OpenStudy (mrnood):

D*pi = C DIVIDE both sides by pi

OpenStudy (sugarmochi):

oo

OpenStudy (sugarmochi):

i get it!

OpenStudy (mrnood):

Well done - post your answer...

OpenStudy (mrnood):

(the question states "to the nearest mm")

OpenStudy (sugarmochi):

im not sure but is it 318mm?

OpenStudy (mrnood):

yes - well done (1000mm/pi)

OpenStudy (sugarmochi):

thanks so much again!! :D